### v2

@@ 3.5.2 Task generation
4 Example system
4.1 System setup
@@
In the following section, we describe an example Mobile Manipulator System based on the <<TaBSASystem>>. It contains its own components based on the <<Scenario>>, <<ScenarioPlugin>>, <<Task>>, <<Robot>>, <<DecAgent>>, <<AgentPlugin>>, and <<EvalFunction>> stereotypes. They are graphically presented in Fig. 7.
Figure 7: Mobile Manipulator System
4.1.1 Tasks
@@
Mobile Manipulator Task is based on the <<Task>> stereotype but implements numerous additional operations and properties (see Fig. 8).
Figure 8: Mobile Manipulator Task
The Fall <<Task>> is meant to imitate the situation in which the <<Robot>> is informed about an elderly falling to the ground. The <<Robot>> moves to the chosen location and waits there for a given time duration. This imitates a brief examination of the elderly health. This <<Task>> has the highest ’priority’, cannot be interrupted, and is terminated about 15 minutes after its ’deadline’. The Patrol <<Task>> resembles the <<Robot>> moving down the path from the starting point to the end. This <<Task>> can be interrupted at any time and cannot die. Patrol’s ’priority’ is determined randomly when it is created, but it can never be higher than the Fall’s ’priority’ value. Pick <<Task>> and Place <<Task>> are not used on their own. They are responsible for the <<Robot>> picking up and placing the object. As our objects are represented merely by the points on the map, the <<Task>>s are performed by standing for a set amount of time near the object. These <<Task>>s received a ’priority’ of 0 for the purposes of the Pick And Place <<Task>> - they aren’t used independently, so they don’t need a realistic ’priority’ value, and we didn’t want their priority to influence complex <<Task>>s.
@@
To represent tasks composed of several simpler tasks, we prepared the Complex <<Task>>. This task possesses a list of ’subtasks’ that are worked on one after another. For this reason, the <<Task>>’s ’estimated_duration’ equals $\sum\limits_{subtasks}(subtask.estimated\_duration)$. The <<Task>>’s ’priority’ was experimentally set to $\max(subtask.priority:subtasks)$ to the sum of priorities of subtasks. The only composite <<Task>> prepared within the system was Pick And Place. It currently consists of 3 subtasks: Pick, Patrol, and Place. This <<Task>> aims to transport an item from one place to another. We set the Pick And Place <<Task>> as not interruptable. As these <<Task>>s utilize objects placed in the <<Environment>>, they tend to start and end close to each other.
4.1.2 Decision agents
We have implemented multiple <<DecAgent>>s to test the effectiveness of different <<Scenario>> execution approaches. Despite their inner workings, all <<DecAgent>>s respect the ’preemptive’ property of the <<Task>>s and won’t interrupt the <<Task>> that shouldn’t be paused during execution.
Simple-longest and Simple-shortest <<DecAgent>>s select the longest and the shortest <<Task>> for the execution, respectively. Additionally, they possess the ’hesitance’ property. Every time the <<Task>> is selected for execution, a random number between 0 and 1 is generated. If it is smaller than the hesitance value, the <<DecAgent>> returns the same <<Task>> as in the previous decision without checking anything (if the <<Task>> is still within the received ’jobs’).
The Distance <<DecAgent>> chooses the <<Task>> that is the furthest from the <<Robot>>. Additionally, the <<DecAgent>> possesses the ratio property used to offset the distance by a fraction of the <<Task>>’s estimated duration. If the ’ratio’ is non-zero, the <<DecAgent>> effectively chooses the <<Task>> with the highest ’dist_score’ calculated by subtracting the ’ratio’ times the <<Task>>’s ’estimated_duration’ from the <<Task>>’s ’distance_from_robot’:
$dist\_score=$ $\max_{jobs}(job.distance\_from\_robot-ratio*$ (2)
$*job.estimated\_duration)$
That means that a higher ratio causes the <<DecAgent>> to favour shorter <<Task>>s.
The Scheduler <<DecAgent>> uses the Request Table <<AgentPlugin>>. Request Table sorts the received <<Task>>s by highest ’priority’. <<Task>>s with the same ’priority’ are then sorted by the earliest ’request_time’. Then the <<AgentPlugin>> creates a schedule of sorts using <<Task>>s’ ’deadlines’ and ’estimated_durations’, returning the lists of scheduled and rejected <<Task>>s. Going through the sorted list, if the <<Task>>’s time from start to finish collides with any other <<Task>> in the scheduled list, the <<Task>> is assigned to the rejected list and vice versa. In the end, the Scheduler <<DecAgent>> receives both of the lists. It returns the <<Task>> from the scheduled list, which time of execution from start to finish contains the current time. If there is no such <<Task>>, the <<DecAgent>> doesn’t return any of them. The <<Task>>s which aren’t worked on at their scheduled time will not be completed in the free time.
The last <<DecAgent>> is called the DQN because it uses the neural network trained using the Deep Q Network [26] (DQN) algorithm to choose the <<Task>> for execution. The network’s input is of the shape [number of <<Task>> types][number of <<Task>>s per type][3] as presented in Fig. 9. The proposed system also contains the interface for the RLlib library [22]. The RLlib contains multiple Reinforcement Learning (RL) state-of-the-art methods that can also be tested for training and benchmarking algorithms scheduling tasks for mobile robots working in dynamic environments, so alternatively, many other RL algorithms can be used in this block. Considering that only Fall, Patrol, and Pick And Place <<Task>>s are directly used within the Service Scenario, in our case, the number of <<Task>> types is 3. For each <<Task>> type, the set number of <<Task>>s is selected by the earliest deadline. Sometimes, the amount of <<Task>> may be too low to fill the whole input. In such cases, the empty places are filled with zeros. For each <<Task>>, three properties are placed inside the network’s input: ’estimated_duration’, ’preemptive’ (represented as integer), and ’distance_from_robot’. <<Task>>s placed within the network’s input are remembered as a list. The network processes the input and outputs the values for each <<Task>>. The highest value marks the position of the <<Task>> to be executed within the saved list. The important process of training the DQN <<DecAgent>> (as per [UC2]) takes place inside its select_task() operation using deep reinforcement learning. After the <<DecAgent>> calculates the new input (shape presented in Fig. 9) it can supply it along with the previous input, selected <<Task>>, and ’reward’ from the <<EvalFunction>> to the training memory used to teach the network. Training the <<DecAgent>> is a long process and requires running multiple <<Scenario>>s using the same <<DecAgent>> (as presented in Fig. 3) as per [UC2.4].
Figure 9: DQNAgent’s network input and output
Apart from the Request Table, our system also contains another <<AgentPlugin>> – Task Predictor. Its purpose is to predict <<Task>>s that can be requested in the near future before they are called. This <<AgentPlugin>> possesses an instance of the Request Table <<AgentPlugin>>. Its most important part is a neural network with Long Short-Term Memory (LSTM) network [31] layers. The plugin is meant to predict <<Task>>s within the set time horizon divided into slots. For each slot, 3 values are calculated: number of <<Task>>s within the slot, number of <<Task>>s set to be executed within the slot (per information from Request Table), and the sum of the ’priorities’ of executable <<Task>>s within the slot. These values, along with the seed used for <<Task>> generation, are passed to the network. The network output consists of the same three values for the same amount of slots but for time horizon counted since one slot in the future. The visualisation of this network’s input and output can be found in Fig. 10. Calculating the differences between the output and the input allows predicting if new <<Task>>s will be added within these time slots in the future.
Figure 10: Task Predictor’s network input and output
4.1.3 Scenario configurations
@@
The example Service Scenario, based on the <<Scenario>> stereotype, has a Mobile Manipulator <<Robot>>, a single <<ScenarioPlugin>> – Duration Estimator Network, and only Mobile Manipulator Task <<Task>>s (Fig. 11). The <<Scenario>> is meant to represent several hours of work for the mobile service <<Robot>>. The <<Robot>> has a set velocity for traveling between <<Task>>s. A certain number of <<Task>>s of each type – Fall, Patrol, and Pick And Place – is generated within these hours. The properties of these <<Task>>s are updated every time by the update_job() operation: estimate_duration() is called, ’distance_from_robot’ is calculated, and every path that needs to be planned within the <<Task>> is refreshed using the current state of the environment.
Figure 11: Service Scenario
Duration Estimator Network <<ScenarioPlugin>> is as the name suggests meant to estimate the duration of the <<Task>>. Despite every <<Task>> in our Mobile Manipulator System having its own estimate, in the real world, these numbers may be offset by an inconsistent environment caused by, for example, humans walking around or misplacement of furniture and objects. Therefore precise prediction of the <<Task>>’s duration is impossible. The plugin consists of a neural network trained offline on data gathered from multiple runs. To mitigate overtraining the network, during data gathering, ’durations’ and ’deadlines’ of <<Task>>s are slightly modified by adding noise and additional <<Task>>s are added randomly to the <<Scenario>>. The network has a single output representing the predicted <<Task>> duration. Its input, on the other hand, consists of the following values: the day when the predicted <<Task>> is performed (seed from <<Task>> configuration), current time, <<Task>>’s ’deadline’, <<Task>>’s current position, <<Task>>’s goal position (if it requires movement, otherwise current position again), <<Task>>’s current ’distance_from_robot’ and <<Task>>’s ’priority’. The structure is presented in Fig. 12.
Figure 12: Duration Estimator Network’s network input and output
4.1.4 Evaluation functions
@@
$\footnotesize reward=\left\{\begin{array}[]{ll}reward\_all\_complete&if\,all\,tasks\,are\\ &\quad completed\\ penalty\_dead\_job&if\,at\,least\,one\,job\\ &\quad is\,dead\\ reward\_job\_complete&if\,job\,was\,just\\ &\quad completed\\ 0&if\,there\,is\,no\,jobs\\ &\quad in\,the\,list\\ reward\_real\_job&if\,real\,job\,was\\ &\quad selected\\ penalty\_nonexistent\_job&if\,nonexistent\,job\\ &\quad was\,chosen\\ \end{array}\right.$ (3)
@@
Statistic Eval aims to numerically assess the work of any <<DecAgent>> it monitors using statistics. This makes it good for every <<DecAgent>> in the <<TaBSASystem>>. It allows the user to verify what the decisions made by the <<DecAgent>> accomplish in the chosen <<Scenario>>. Statistic Eval returns the ’terminate’ flag if all <<Task>>s are completed, one of the <<Task>>s is terminated or the <<DecAgent>> switches between <<Task>>s and returns to the same one for the third time within 3 minutes. Other useful values returned by the evaluation function on each iteration are represented by the following properties:
  • full_travel_distance – the full travel distance of the <<Robot>> calculated by comparing the <<Robot>>’s current pose to the remembered ’last_robot_pos’,
  • num_of_tasks_completed – the number of completed <<Task>>s of every type in ’task_types’,
  • task_completion_to_deadline – the difference between the <<Task>>’s completion time and the original ’deadline’,
  • task_completion_to_deathtime – the difference between the <<Task>>’s completion time and potential time of termination, None if a <<Task>> can’t be terminated,
  • num_of_tasks_interrupted – number of <<Task>>s’ interruptions per <<Task>> type in ’task_types’,
  • task_interruptions – number of <<Task>>s’ interruptions for each <<Task>> in <<Scenario>>’s ’task’ list,
  • num_of_human_abandonement – number of <<Task>>s’ interruptions for each <<Task>> and number of times the <<Robot>> has abandoned a human in need.
The last statistic is specific to our system and represents the number of situations where the <<Robot>> drove within ’abandonment_distance’ of any Fall <<Task>> without working on it.
4.2 Example benchmarking sessions
4.2.1 Decision agent training
As mentioned before, DQN <<DecAgent>> requires a trained neural network to operate. Training this network may be considered an example of a successful benchmarking session utilising our Mobile Manipulator System. The <<DecAgent>> was taught using reinforcement learning (DQN algorithm). The training consisted of the <<DecAgent>> rehearsing numerous <<Scenario>>s. During the training, the <<Scenario>>s reflected 4 hours of social <<Robot>> work, during which the <<DecAgent>> was expected to complete 12 <<Task>>s of each of the 3 types. In each, the <<Task>>s were generated using a new seed. The DQN <<DecAgent>>’s actions were evaluated using DQN Eval, and the reward received during this stage was used as a reward for reinforcement learning. Learning took several million steps, each representing 5 seconds of virtual time. Individual <<Scenario>>s ended if all <<Task>>s were completed, one of the <<Task>>s died or <<Task>> completion exceeded the <<Scenario>> time. Once training was complete, the network was saved for further use. During the project, the training was repeated several times to verify differences in the performance of the different network structures and the value of the reward for individual actions.
4.2.2 Agent comparison
@@
Another example of a benchmarking session could be an attempt to compare the performance of individual decision-making <<DecAgent>>s. Such an exercise makes it possible to determine the best one and to identify potential errors in its performance. Initially, 50 <<Scenario>>s with random <<Task>>s were generated for each <<DecAgent>> under test. The <<Scenario>>s lasted 4 hours and contained 12 <<Task>>s of each of the three types. This time the Statistic Eval was used to evaluate the performance of the <<DecAgent>>s. The results of each run were saved for analysis. Some of the results for six different <<DecAgent>>s (Distance with ’ratio’ 0.5, two versions of DQN, Scheduler, Simple-longestwith ’hesitance’ 0.5, and Simple-shortest with ’hesitance’ 0.5) are shown below, as they presented us with useful information about the <<DecAgent>>s and allowed us to improve them.
@@
Fig. 13 includes statistics for scenario termination cause for different types of <<DecAgent>>s presented as bars. We can see that Simple-shortest and Distance <<DecAgent>>s were moderately successful in completing more than half of the <<Scenario>>s. Scheduler and Simple-longest <<DecAgent>>s both failed for different reasons. If two Fall <<Task>>s were to be scheduled at the same time, Scheduler would only complete one of them. Simple-longest agent, on the other hand, prioritizes long <<Task>>s, which causes him to start with Pick And Place <<Task>>s, which tend to be longer than the others. DQN <<DecAgent>> didn’t have a good outcome during this test either, but something else was also spotted. After training the first DQN <<DecAgent>>, a lot of <<Scenario>>s ended due to running out of time. This fact prompted us to reevaluate the code and discover that the penalty for doing nothing was indeed wrongly applied as a positive number. This was fixed for the second network, which, as one can see, doesn’t display the same behaviour.
@@
Fig. 14 shows the absolute difference between <<Task>>s’ completion times and their original ’deadlines’ sorted by <<Task>> type. As can be seen, one of the <<DecAgent>>s seems to be much better than the others - Scheduler - as it displayed much smaller disparities. This means that <<Task>>s worked on by this agent were completed close to the original ’deadlines’, which may be a desirable behaviour in some cases. All of the other <<DecAgent>>s start performing <<Task>>s as fast as they come, which causes the differences to be quite large. We can also observe that the biggest differences are visible for Pick And Place <<Task>>s, which suggests that they are usually the longest to complete or are susceptible to moving pedestrians.
@@
Figure 14: Difference between real completion time and the ’deadline’ for different <<Task>> types for different <<DecAgent>>s
@@
Fig. 15 shows the number of <<Task>>s of each type completed by each <<DecAgent>> during the mentioned 50 <<Scenario>>. These plots also confirm that most <<DecAgent>>s work as intended despite failing the <<DecAgent>>s. For example, the Scheduler has more Fall <<Task>>s completed than other types, as they have the highest manual ’priority’. On the other hand, Simple-longest chooses to work on the longest <<Task>>s first, so it prioritizes Pick And Place <<Task>>s. The biggest surprise is the first DQN, which seems to have completed almost as many fall <<Task>>s as the Distance <<DecAgent>> but didn’t complete any <<Scenario>>s. This further confirms the need to retrain the <<DecAgent>> with different parameters.
Figure 15: Number of completed <<Task>>s per type for different <<DecAgent>>s
@@
After that, another experiment was performed, in which each <<DecAgent>> completed the same 100 <<Scenario>>s (same seed used for <<Task>> generation). This allowed us to inspect the differences in <<DecAgent>>s’ behaviours closely. Figures 16 and 17 show how the same scenario was completed by Simple-shortest and Distance <<DecAgent>>s. The horizontal axis represents the passage of time, while the currently performed <<Task>> is identified by its colour and order in the <<Task>> list of the <<Scenario>>. For example, a green line at 6 represents the sixth generated Pick And Place task. Comparing the plots, we can see that the <<Task>>s were completed in a slightly different order at the beginning due to differences in decision-making. On the other hand, since about 1700th second, the outcomes are almost identical, probably because these <<Task>>s were given to the agents at the same time late into the day.
Figure 16: Scenario course as performed by Simple-shortest <<DecAgent>>
Figure 17: Scenario course as performed by Distance <<DecAgent>>
5 Summary
This article tackles the problem of assessing task-scheduling algorithms in the mobile robot domain. We propose a benchmarking framework that encapsulates the scheduling algorithm with the <<DecAgent>> interface. Numerous algorithms are optimised for given robot applications, robot types, or environments. Therefore, our framework allows the configuration of the robot’s kinematics, tasks, environment, and testing scenarios. By the configuration, the framework reflects the desired robot system and its environment setup. The system is open-source, and its repository is shared [^2] https://github.com/RCPRG-ros-pkg/Smit-Sim. Therefore, our experiments can be easily re-executed, and the framework can be used to evaluate new or well-known algorithms in the domain of mobile robots.
The framework incorporates the generators for schedules of <<Task>> requests, trajectories of moving humans, and maps with objects for manipulation and rooms. These generators can be used in training task-scheduling algorithms and for statistic-based evaluation of these or other algorithms.
We have validated the framework by configuring it for a mobile service <<Robot>> application, teaching AI-based algorithms using DQN, and assessing it against basic scheduling algorithms. As a result of the framework application, we could identify flaws in both coding and decision-making. Ultimately, we could understand and graphically present the core differences between the selected algorithms at chosen configurations. In the given test case, scheduling algorithms that chose the shortest (Simple-shortest <<DecAgent>> with ’hesitance’ 0.5) or the furthest (Distance <<DecAgent>> with ’ratio’ 0.5) task managed to complete the <<Scenario>>s respectively  90% and  60% of the time. They also didn’t seem to favour any specific type of <<Task>>. The algorithm choosing the longest <<Task>> (Simple-longest<<DecAgent>> with ’hesitance’ 0.5) did the worst (with all <<Scenario>>s fatally failed), followed closely by the AI-based ones (both versions of DQN <<DecAgent>>) and the one based on the strict schedule (Scheduler <<DecAgent>>). In the first case, the <<DecAgent>> favours executing Pick And Place, while the others Fall <<Task>>s. Of all the algorithms, only the schedule-based one completed the <<Task>>s almost at the set ’deadline’.
From testing the algorithms in the proposed benchmarking system, we not only select the best algorithm but also learn about the <<Robot>> application. During the investigation of our example, we learned that both of the best algorithms managed to get ahead by completing as many <<Task>>s as fast as possible by favouring shorter <<Task>>s. Additionally, the Distance <<DecAgent>> avoided tasks that are close to each other. Pick And Place <<Task>>s tend to start and end close to each other and are often the longest to perform, so algorithms that favour them for any reason are likely to fail. Seeing that most scenarios ended due to the "dead" status (which happens when Fall <<Task>> is not completed for some time after its ’deadline’), we find that keeping deadlines is also critical in this application. Thus, the scheduling algorithms that are aware of deadlines should perform better than the ones that are not. On the other hand, a <<DecAgent>> may concentrate on Fall <<Task>>s but can still fail if it uses a subpar scheduling algorithm.
Scheduling tasks of mobile robots is challenging in a dynamic environment. However, based on the experience gained in this research, we find the advanced scheduling of manipulation tasks the next valuable target. Advanced scheduling must involve semantic planning of the tasks, which is a state-of-the-art solution to managing complexity in robot tasks [29].
[…]

### v3

@@ 2.4.2 Task generation
2.5 Example benchmarking sessions
2.5.1 Decision agent training
As mentioned before, DQN <<DecAgent>> requires a trained neural network to operate. Training this network may be considered an example of a successful benchmarking session utilising our Mobile Manipulator System. The <<DecAgent>> was taught using reinforcement learning (DQN algorithm). The training consisted of the <<DecAgent>> rehearsing numerous <<Scenario>>s. During the training, the <<Scenario>>s reflected 4 hours of social <<Robot>> work, during which the <<DecAgent>> was expected to complete 12 <<Task>>s of each of the 3 types. In each, the <<Task>>s were generated using a new seed. The DQN <<DecAgent>>’s actions were evaluated using DQN Eval, and the reward received during this stage was used as a reward for reinforcement learning. Learning took several million steps, each representing 5 seconds of virtual time. Individual <<Scenario>>s ended if all <<Task>>s were completed, one of the <<Task>>s died or <<Task>> completion exceeded the <<Scenario>> time. Once training was complete, the network was saved for further use. The training was repeated several times to verify differences in the performance of the different network structures and the value of the reward for individual actions.
3 Illustrative example
@@
In the following section, we describe an example Mobile Manipulator System based on the <<TaBSASystem>>. It contains its own components based on the <<Scenario>>, <<ScenarioPlugin>>, <<Task>>, <<Robot>>, <<DecAgent>>, <<AgentPlugin>>, and <<EvalFunction>> stereotypes. They are graphically presented in Fig. 5(a).
(a) Mobile Manipulator System
(b) Mobile Manipulator Task
Figure 5: Mobile Manipulator system definition
3.1 Tasks
@@
Mobile Manipulator Task is based on the <<Task>> stereotype but implements numerous additional operations and properties (see Fig. 5(b)).
@@
To represent tasks composed of several simpler tasks, we prepared the Complex <<Task>>. This task possesses a list of ’subtasks’ that are worked on one after another. For this reason, the <<Task>>’s ’estimated_duration’ equals $\sum\limits_{subtasks}(subtask.estimated\_duration)$. The <<Task>>’s ’priority’ was experimentally set to $\max(subtask.priority:subtasks)$ to the sum of priorities of subtasks. The only composite <<Task>> prepared within the system was Pick And Place. This <<Task>> aims to transport an item from one place to another.
3.2 Decision Agents
  • 1.
    Simple-longest and Simple-shortest <<DecAgent>>s select the longest and the shortest <<Task>> for the execution, respectively.
  • 2.
    The Distance <<DecAgent>> chooses the <<Task>> that is the furthest from the <<Robot>>.
  • 3.
    The Scheduler <<DecAgent>> uses the Request Table <<AgentPlugin>>. Request Table sorts the received <<Task>>s by highest ’priority’. <<Task>>s with the same ’priority’ are then sorted by the earliest ’request_time’.
  • 4.
    DQN <<DecAgent>> implements the neural network trained using the Deep Q Network [9] (DQN) algorithm to choose the <<Task>> for execution.
3.3 Scenario
@@
The example Service Scenario, based on the <<Scenario>> stereotype, has a Mobile Manipulator <<Robot>>, a single <<ScenarioPlugin>> – Duration Estimator Network, and only Mobile Manipulator Task <<Task>>s. The <<Scenario>> is meant to represent several hours of work for the mobile service <<Robot>>. The <<Robot>> has a set velocity for traveling between <<Task>>s. A certain number of <<Task>>s of each type – Fall, Patrol, and Pick And Place – is generated within these hours. The properties of these <<Task>>s are updated every time by the update_job() operation: estimate_duration() is called, ’distance_from_robot’ is calculated, and every path that needs to be planned within the <<Task>> is refreshed using the current state of the environment.
3.4 Evaluation functions
@@
[…]
@@
Statistic Eval aims to numerically assess the work of any <<DecAgent>> it monitors using statistics. This makes it good for every <<DecAgent>> in the <<TaBSASystem>>. It allows the user to verify what the decisions made by the <<DecAgent>> accomplish in the chosen <<Scenario>>. Statistic Eval returns the ’terminate’ flag if all <<Task>>s are completed, one of the <<Task>>s is terminated or the <<DecAgent>> switches between <<Task>>s and returns to the same one for the third time within 3 minutes. In the documentation, there are other useful values returned by the evaluation function on each iteration.
3.5 Experimental results
@@
Initially, 50 <<Scenario>>s with random <<Task>>s were generated for each <<DecAgent>> under test. The <<Scenario>>s lasted 4 hours and contained 12 <<Task>>s of each of the three types. This time the Statistic Eval was used to evaluate the performance of the <<DecAgent>>s. The results of each run were saved for analysis. Some of the results for six different <<DecAgent>>s (Distance with ’ratio’ 0.5, two versions of DQN, Scheduler, Simple-longestwith ’hesitance’ 0.5, and Simple-shortest with ’hesitance’ 0.5) are shown below, as they presented us with useful information about the <<DecAgent>>s and allowed us to improve them.
@@
Fig. 6 includes statistics for scenario termination cause for different types of <<DecAgent>>s presented as bars. We can see that Simple-shortest and Distance <<DecAgent>>s were moderately successful in completing more than half of the <<Scenario>>s. Scheduler and Simple-longest <<DecAgent>>s both failed for different reasons. During the framework execution, we identified unwanted behaviour of the first DQN <<DecAgent>>. It has terminated a lot of <<Scenario>>s due to running out of time. This fact prompted us to reevaluate the code and discover that the penalty for doing nothing was indeed wrongly applied as a positive number. This was fixed for the second network, which, as one can see, doesn’t display the same behaviour.
@@
Fig. 7(a) shows the absolute difference between <<Task>>s’ completion times and their original ’deadlines’ sorted by <<Task>> type. As can be seen, one of the <<DecAgent>>s seems to be much better than the others - Scheduler - as it displayed much smaller disparities. This means that <<Task>>s worked on by this agent were completed close to the original ’deadlines’, which may be a desirable behaviour in some cases.
@@
(a) Difference between real completion time and the ’deadline’ for different <<Task>> types for different <<DecAgent>>s
(b) Number of completed <<Task>>s per type for different <<DecAgent>>s
Figure 7: Comparison of the scheduling algorithms
@@
Fig. 7(b) shows the number of <<Task>>s of each type completed by each <<DecAgent>> during the 50 <<Scenario>>s. These plots also confirm that most <<DecAgent>>s work as intended despite failing the <<DecAgent>>s. For example, the Scheduler has more Fall <<Task>>s completed than other types, as they have the highest manual ’priority’. On the other hand, Simple-longest chooses to work on the longest <<Task>>s first, so it prioritizes Pick And Place <<Task>>s. The biggest surprise is the first DQN, which seems to have completed almost as many fall <<Task>>s as the Distance <<DecAgent>> but didn’t complete any <<Scenario>>s. While the DQN-based decision agent performed poorly in our experiments, this outcome highlights a central feature of the proposed framework: its ability to reveal limitations of specific approaches under controlled, reproducible conditions. The framework is designed precisely for such evaluations, enabling algorithm developers to identify, analyze, and address weaknesses in their methods. Thus, the weaker performance of DQN in this setting demonstrates the effectiveness of the proposed framework in revealing algorithmic limitations or providing valuable insights for further improvements.
@@
After that, another experiment was performed, in which each <<DecAgent>> completed the same 100 <<Scenario>>s (same seed used for <<Task>> generation). This allowed us to inspect the differences in <<DecAgent>>s’ behaviours closely. Figures 8(a) and 8(b) show how the same scenario was completed by Simple-shortest and Distance <<DecAgent>>s. The horizontal axis represents the passage of time, while the currently performed <<Task>> is identified by its colour and order in the <<Task>> list of the <<Scenario>>. For example, a green line at 6 represents the sixth generated Pick And Place task. Comparing the plots, we can see that the <<Task>>s were completed in a slightly different order at the beginning due to differences in decision-making. On the other hand, since about 1700th second, the outcomes are almost identical, probably because these <<Task>>s were given to the agents at the same time late into the day.
(a) Scenario course as performed by Simple-shortest <<DecAgent>>
(b) Scenario course as performed by Distance <<DecAgent>>
Figure 8: Comparison of the example scenario execution
4 Impact
The presented framework has demonstrated significant utility in the quantitative evaluation of scheduling algorithms, referred to as Decision Agents, for service robots operating in dynamic environments. By executing each agent within identical, simulated scenarios that include randomized human movement and diverse tasks, the framework enables fair, reproducible comparisons. It supports both broad benchmarking across general conditions and targeted testing in application-specific setups. The experiments confirmed the framework’s capability to reveal subtle behavioral differences between agents, highlight performance patterns, and, crucially, expose implementation or configuration errors—such as misapplied penalties in one of the tested DQN agents. Through detailed statistical outputs and visual comparisons, users can assess algorithm versatility or select the most suitable scheduling strategy for a given robotic application, thereby accelerating both algorithm development and deployment-readiness.
5 Conclusions
We proposed and validated a practical benchmarking framework for evaluating scheduling algorithms in mobile service robotic applications. The framework enables configurable simulation of environments, tasks, robot kinematics, and human activity, allowing reproducible and comparative testing of decision-making algorithms under identical conditions. The experiments demonstrated the framework’s effectiveness in identifying both behavioral characteristics and implementation flaws, highlighting its diagnostic value. By visualizing differences in task execution, completion timing, and failure causes, the framework supports both algorithm tuning and application-driven selection. The presented results confirm that the framework can effectively benchmark diverse decision-making agents, from heuristic schedulers to learning-based approaches. The fact that certain algorithms, such as DQN, underperformed demonstrates the framework’s usefulness. It provides a fair and transparent way to assess algorithms and to identify cases where they can be further refined.
It is open-source and extensible [^2] https://github.com/RCPRG-ros-pkg/Smit-Sim, therefore it serves as a ready-to-use tool for practitioners aiming to improve scheduling strategies or adapt them to specific robot deployments. Finally, the insights gained suggest that task-deadline awareness and preference for shorter tasks increase scheduling robustness in realistic scenarios, while future work should target semantic-level planning for more complex robot activities.